# A look on equations describing pseudospherical surfaces

### v1

@@ Abstract
We revisit the notion of equations describing pseudospherical surfaces starting from its roots influenced by the AKNS system until current research topics in the field.
Keywords Equations describing pseudospherical surfaces  $\cdot$ Camassa-Holm equation
Dedicatory: To Professor Keti Tenenblat on occasion of her 80th birthday.
@@ 1 Introduction
In [1] the authors considered the following linear system
$\partial_{x}\begin{pmatrix}v_{1}\\ \\ v_{2}\end{pmatrix}=\begin{pmatrix}-i\zeta&q\\ \\ r&i\zeta\end{pmatrix}\begin{pmatrix}v_{1}\\ \\ v_{2}\end{pmatrix},$ (1.0.1)
$\partial_{t}\begin{pmatrix}v_{1}\\ \\ v_{2}\end{pmatrix}=\begin{pmatrix}A&B\\ \\ C&D\end{pmatrix}\begin{pmatrix}v_{1}\\ \\ v_{2}\end{pmatrix},$ (1.0.2)
where, $q$ and $r$ are functions of $(x,t)$ whereas $A$, $B$, $C$ and $D$ depend on $x,\,t$ and $\zeta$, which may eventually depend on t. The field variables $x$ and $t$ can be regarded as space and time, respectively.
Equation (1.0.1) is a scattering problem while (1.0.2) is an evolution problem.
From the cross differentiation and the integrability condition
$\partial_{t}\partial_{x}\begin{pmatrix}v_{1}\\ \\ v_{2}\end{pmatrix}=\partial_{x}\partial_{t}\begin{pmatrix}v_{1}\\ \\ v_{2}\end{pmatrix}$
(1.0.1) and (1.0.2) give
$\begin{pmatrix}\partial_{x}A&\partial_{x}B\\ \\ \partial_{x}C&\partial_{x}D\end{pmatrix}-\begin{pmatrix}0&\partial_{t}q\\ \\ \partial_{t}r&0\end{pmatrix}=i\partial_{t}\zeta\begin{pmatrix}-1&0\\ \\ 0&1\end{pmatrix}+\begin{pmatrix}qC-rB&q(D-A)-2i\zeta B\\ \\ r(A-D)+2i\zeta C&rB-qC\end{pmatrix}.$
The $2\times 2$ matrix in (1.0.1) is traceless, as are all but the first in the above equation. If we impose that the second matrix in (1.0.2) be traceless, then all matrices we have so far will share this property. Without loss of generality, we may assume $D=-A$.
@@
$\left\{\begin{array}[]{lcl}\partial_{x}A&=&qC-rB,\\ \\ \partial_{x}B+2i\zeta B&=&\partial_{t}q-2Aq,\\ \\ \partial_{x}C-2i\zeta C&=&\partial_{t}r+2Ar.\end{array}\right.$ (1.0.3)
The reader may be wondering: What is the relevance of (1.0.3)? This is a very good question.
In [1] it was shown that a number of very important models, such as the KdV and the sine-Gordon equations, can be obtained from (1.0.3) depending on the choices of the involved functions. Moreover, equations arising from the compatibility condition (1.0.3) are said to be integrable (à la AKNS).
Let us give two examples illustrating how equations can be obtained from (1.0.3). The first one is a very classical case, while the second one is focused on a shallow water model unknown ty the time [1] was published.
Example 1.1.
@@
Then (1.0.3) is equivalent to the sine-Gordon (sG) equation
$u_{tx}=\sin{u}.$ (1.0.4)
It is worth mentioning that the sG equation was the first PSS equation to be discovered, even before the concept was brought to light. This comes from the fact that the sG plays vital a vital role in the context of the geometry of surfaces with constant and negative curvatures [36, Chapter 1].
Example 1.2.
Consider
$\begin{array}[]{l}A=\displaystyle{\frac{1}{2}\Big{(}u_{x}-\eta u-\frac{\beta}{\eta}\Big{)},\,\,\,B=\frac{1}{2}\Big{(}\beta u-1+\frac{\beta u}{\eta^{2}}+u+\frac{u}{\eta^{2}}\Big{)}}\\ \\ C=\displaystyle{\frac{1}{2}\Big{(}\frac{2}{\eta^{2}}-2um+\beta u+\frac{\beta u}{\eta^{2}}-u-2\frac{\beta u_{x}}{\eta}+2\frac{u_{x}}{\eta}-u-\frac{2\beta+u}{\eta^{2}}\Big{)}},\\ \\ r=\displaystyle{\frac{1}{2}\Big{(}1-2m-\beta+\frac{\beta+1}{\eta^{2}}\Big{)},\,\,q=\frac{1}{2}\Big{(}\frac{\beta-1}{\eta^{2}}-1-\beta\Big{)},\,\,m=u-u_{xx},\,\,\eta=-2i\zeta},\end{array}$
where the parameters satisfy the constraint
$\eta^{4}-\eta^{2}+\beta^{2}\eta^{2}=\beta^{2}+1-2\beta.$
Then (1.0.3) is equivalent to the Camassa-Holm (CH) equation
$u_{t}-u_{txx}+3uu_{x}=2u_{x}u_{xx}+uu_{xxx},$ (1.0.5)
or its equivalent form
$m_{t}+2um_{x}+u_{x}m=0.$ (1.0.6)
Let us now introduce a geometric view on (1.0.1)–(1.0.2). By $\{dx,dt\}$ we denote the usual dual basis of $\mathbb{R}^{2}$. Consider eight one-forms $\Omega_{ij},\,\Psi_{ij}$, $1\leq i,j\leq 2$ and define one-form matrices $\Omega=(\Omega_{ij})$, $\Psi=(\Psi_{ij})$. If $d$ denotes the usual exterior derivative, then we have
$d\Omega=(d\Omega_{ij})$
@@
$\Omega=\begin{pmatrix}-i\zeta dx+Adt&qdx+Bdt\\ \\ rdx+Cdt&i\zeta dx-Adt\end{pmatrix}$ (1.0.7)
@@
$dv=\Omega v,$ (1.0.8)
where $v=(v_{1}\quad v_{2})^{t}$ (the superscript $t$ means the usual matrix transposition). From (1.0.6) we obtain
@@
$d\Omega=\Omega\wedge\Omega.$ (1.0.9)
@@
then (1.0.7) gives the equation
$\partial_{t}X+\partial_{x}T+[X,T]=0.$ (1.0.10)
Equation (1.0.8) is better known as the zero curvature representation, and it reveals to us that (1.0.1)–(1.0.2) can be seen as a geometric object.
Equations (1.0.7) are a different, but equivalent, way to write (1.0.1)–(1.0.2). whereas (1.0.9) implies to (1.0.3).
The remainder of the paper is as follows: in Section 2 we revisit the main aspects of the works by Sasaki [38] and Chern and Tenenblat [8], connecting equations obtained from the AKNS method and the geometry of abstract 2-manifolds with constant and negative Gaussian curvature. In Section 3 we discuss some recent results from the author [16, 19, 20, 21, 37] that connect the theory developed by Chern and Tenenblat with problems concerning not only equations, but equations subject to some condition (initial data). In some parts, the presentation may closely follow the original references.
@@ 2 Equations describing pseudospherical surfaces
Chern and Tenenblat pioneered the concept of pseudospherical equations [8], establishing a connection between certain partial differential equations and smooth two-dimensional Riemannian manifolds. Their contribution gained particular significance following a notable insight by Sasaki, who had earlier demonstrated that solutions to specific integrable equations—those arising from the AKNS formalism [1]—are intimately related to the intrinsic geometry of surfaces with constant negative Gaussian curvature ${\cal K}=-1$ [38].
@@
A pseudospherical surface (PSS) is a two-dimensional Riemannian manifold whose Gaussian curvature is constant and negative.
These foundational results laid the groundwork for a new intersection between surface theory and the analysis of PDEs. Since then, this perspective has shaped a substantial body of literature, including [2, 5, 6, 7, 16, 19, 31, 32, 33, 34, 35, 37], among many others. A more comprehensive introduction to the topic may be found in [39, Chapter 1].
By definition, a pseudospherical surface (abbreviated as PSS, also used in the plural) is a two-dimensional Riemannian manifold whose Gaussian curvature remains constant and negative. According to Hilbert’s theorem [15, Section 5-11], such manifolds cannot be isometrically immersed in $\mathbb{R}^{3}$ as complete surfaces, due to their negative curvature—see also [28, page 439] and [27]. For this reason, they are frequently referred to as abstract surfaces in the literature.
${\cal E}(x,t,u,u_{(1)},\cdots,u_{(n)})=0.$ (2.0.1)
@@
A differential equation (2.0.1) is said to describe a pseudospherical surface, or it is said to be of pseudospherical type, if it is a necessary and sufficient condition for the existence of differentiable functions $f_{ij}$, $1\leq i,j\leq 3$, such that the forms
$\omega_{i}=f_{i1}dx+f_{i2}dt,\quad 1\leq i\leq 3,$ (2.0.2)
satisfy the structure equations of a pseudospherical surface
$d\omega_{1}=\omega_{3}\wedge\omega_{2},\quad d\omega_{2}=\omega_{1}\wedge\omega_{3},\quad d\omega_{3}=\omega_{1}\wedge\omega_{2}.$ (2.0.3)
Sasaki noticed in his work [38] that by defining one-forms $\omega_{1}$, $\omega_{2}$ and $\omega_{3}$ suitably from (1.0.1) and (1.0.2), one could obtain the one-form matrix
$\Omega=\frac{1}{2}\begin{pmatrix}\omega_{2}&\omega_{1}-\omega_{3}\\ \omega_{1}-\omega_{3}&-\omega_{2}\end{pmatrix},$ (2.0.4)
$\Sigma:=d\Omega-\Omega\wedge\Omega.$ (2.0.5)
Then $\Sigma=0$ if and only if (2.0.3) holds. As a result, if (2.0.1) is the compatibility condition of (1.0.3), then its solutions satisfy (2.0.3), that is, the structure equations for a two-dimensional Riemannian manifold with first fundamental form, or metric, given by
$I=\omega_{1}^{2}+\omega_{2}^{2}.$ (2.0.6)
In practical terms, given a triad of one-forms $\omega_{1}$, $\omega_{2}$ and $\omega_{3}$, and an equation (2.0.1), we can check if they describe a PSS surface in the following way: let us define a matrix of one-forms $\Omega$ by (2.0.4). If, when restricted to the manifold determined by the solutions of (2.0.1), the matrix $\Sigma$ vanishes, we say that (2.0.1) is a PSS equation, and the triad $\{\omega_{1},\,\omega_{2},\,\omega_{3}\}$ satisfies the structure equations (2.0.3) of a PSS with Gaussian curvature ${\cal K}=-1$. As a result, any open, simply connected set $U$ where the one-forms $\omega_{1}$ and $\omega_{2}$ are defined and $\omega_{1}\wedge\omega_{2}\neq 0$ everywhere, is endowed with a PSS structure. We can naively see this condition as a requirement for $\{\omega_{1},\omega_{2}\}$ to be a linearly independent set and thus be the basis of a plane.
Definition 2.3.
Assume that (2.0.1) is an equation describing a PSS and let $u$ be a solution of it. We say that $u$ is a generic solution if the one-forms (2.0.2) satisfy the condition $\omega_{1}\wedge\omega_{2}\neq 0$.
@@
$\left\{\begin{array}[]{lcl}\omega_{1}&=&\displaystyle{\frac{1}{\eta}\sin{u}dt,}\\ \\ \omega_{2}&=&\displaystyle{\eta dx+\frac{1}{\eta}\cos{u}dt},\\ \\ \omega_{3}&=&u_{x}dx.\end{array}\right.$ (2.0.7)
@@
$\left\{\begin{array}[]{lcl}d\omega_{1}-\omega_{3}\wedge\omega_{2}&\equiv&\displaystyle{0}\\ \\ d\omega_{2}-\omega_{1}\wedge\omega_{3}&\equiv&\displaystyle{0},\\ \\ d\omega_{3}-\omega_{1}\wedge\omega_{2}&=&-(u_{tx}-\sin{u})dx\wedge dt.\end{array}\right.$
As a result, on the solutions of the sG equation (1.0.4), the forms (2.0.7) satisfy (2.0.3). Since
@@
then $u$ will be a generic solution for (1.0.4) provided that $u\notin\{k\pi,\,\,\,k\in\mathbb{Z}\}$.
@@
$\begin{array}[]{lcl}\omega_{1}&=&\displaystyle{\Big{(}\frac{\lambda}{2}+\frac{1}{2\lambda}-m\Big{)}dx+\Big{(}um+\frac{\lambda}{2}u-\frac{u}{2\lambda}-\frac{1}{2}-\frac{\lambda^{2}}{2}\Big{)}dt},\\ \\ \omega_{2}&=&-u_{x}dt,\\ \\ \omega_{3}&=&\displaystyle{\Big{(}m+\frac{1}{2\lambda}-\frac{\lambda}{2}\Big{)}dx+\Big{(}\frac{\lambda^{2}}{2}-\frac{1}{2}-\frac{u}{2\lambda}-\frac{\lambda}{2}u-um\Big{)}dt}.\end{array}$ (2.0.8)
@@
$\begin{array}[]{lcl}d\omega_{1}-\omega_{3}\wedge\omega_{2}&=&\Big{(}u_{t}-u_{txx}+3uu_{x}-2u_{x}u_{xx}-uu_{xxx}\Big{)}dx\wedge dt,\\ \\ d\omega_{2}-\omega_{1}\wedge\omega_{3}&=&0,\\ \\ d\omega_{3}-\omega_{1}\wedge\omega_{2}&=&-\Big{(}u_{t}-u_{txx}+3uu_{x}-2u_{x}u_{xx}-uu_{xxx}\Big{)}dx\wedge dt,\end{array}$ (2.0.9)
meaning that on the generic solutions of the CH equation (1.0.5), (2.0.8) define a two-dimensional Riemannian manifold with first fundamental form
$\begin{array}[]{lcl}I&=&\displaystyle{\Big{[}m^{2}-\Big{(}\lambda+\frac{1}{\lambda}\Big{)}m+\frac{1}{4}\Big{(}\lambda+\frac{1}{\lambda}\Big{)}^{2}\Big{]}dx^{2}}+\\ \\ &&\displaystyle{\Big{[}-um^{2}+\frac{1}{\lambda}um+\frac{1}{4}\Big{(}\lambda^{2}-\frac{1}{\lambda^{2}}\Big{)}u+\lambda\Big{(}\frac{\lambda}{2}+\frac{1}{2\lambda}\Big{)}m-\lambda\Big{(}\frac{\lambda}{2}+\frac{1}{2\lambda}\Big{)}^{2}\Big{]}dxdt}+\\ \\ &&\displaystyle{\Big{[}um^{2}+u^{2}m\Big{(}\lambda-\frac{1}{\lambda}\Big{)}+u_{x}^{2}-\lambda\Big{(}\lambda+\frac{1}{\lambda}\Big{)}um+\frac{1}{4}\Big{(}\lambda+\frac{1}{\lambda}\Big{)}^{2}u^{2}-\frac{\lambda}{2}\Big{(}\lambda+\frac{1}{\lambda}\Big{)}^{2}}\\ \\ &&\displaystyle{+\frac{\lambda^{2}}{4}\Big{(}\lambda+\frac{1}{\lambda}\Big{)}^{2}\Big{]}dt^{2}}.\end{array}$
@@
$m=\frac{\lambda}{2}+\frac{1}{2\lambda}$ (2.0.10)
@@
Remark 2.1.
It is possible that a PSS equation possesses more than one representation of the forms satisfying the structure equations (2.0.3). If we substitute the choices made in Example 1.2 into (1.0.7) will lead to different one-forms (see [32, Theorem 1]) than those in (2.0.8) for the CH equation, although (1.0.10) would be equivalent to the CH equation (1.0.5). However, any of the triads given either in [32, Theorem 1] or Example 1.2 are enough to ensure that the CH equation describes PSS.
Remark 2.2.

### v4

@@ Abstract
We revisit the notion of equations describing pseudospherical surfaces, starting from the works by Sasaki, whose roots were influenced by the AKNS system, the works by Chern and Tenenblat, until current research topics in the field relating to Cauchy problems involving these equations and their geometric consequences.
keywords:
Equations describing pseudospherical surfaces, sine-Gordon equation, Camassa-Holm equation
[…]
Acknowledgements.
I would like to thank Professor Tiago Macedo for his invitation to contribute to CM with this review paper. I am also thankful to CNPq (grant n. 310074/2021-5) and FAPESP (grant n. 2024/01437-8) for financial support. I am also thankful to the referees for their comments and suggestions that have improved an earlier version of this manuscript. This article is dedicated to Professor Keti Tenenblat on the occasion of her 80th birthday.
@@ 1 Introduction
The deep relations between partial differential equations (PDEs) and geometry date back to the XIX century, when a non-linear partial differential equation was shown to be related to the Gaussian curvature of a surface. Since then, a number of non-linear equations have been proved to be of crucial relevance to better understand geometric aspects of surfaces.
While in the XIX century the connections between PDEs and geometry was established, in the course of the XX century we witnessed a deep relation between equations coming from mathematical physics, notably from hydrodynamics, and their connections with the geometry of surfaces with constant Gaussian curvature.
After the WW II, the KdV equation was rediscovered (for some historical information, see the books by Drazin and Johnson [13] or Kasman [25]), and during the 1960’s it received considerable attention due to the discovery of the soliton by Zabusky and Kruskal [42], and the techniques proposed by Gardner et. al. for solving it [20]. These are two of the most prominent [^1] In order to infer the relevance of [42] and [20], according to Scopus, they have received 3,340 and 3,706 citations, respectively. Data checked on December 12, 2025. works in the field of integrable systems developed in the 60s. After these works the KdV equation attracted spotlights and its relevance was reinforced by a series of works published in the Journal of Mathematical Physics by Gardner, Miura, Kruskal and co-workers, see [28] and references therein. See also [13, 25].
The works [20, 28, 42] were rather influential in the field of non-linear science during the late 60s and early 70s, leading to the development of a new method for finding equations sharing the nice properties of the KdV equation: in [1] the authors considered the following linear system
$\partial_{x}\begin{pmatrix}v_{1}\\ \\ v_{2}\end{pmatrix}=\begin{pmatrix}-i\zeta&q\\ \\ r&i\zeta\end{pmatrix}\begin{pmatrix}v_{1}\\ \\ v_{2}\end{pmatrix}$, (1)
$\partial_{t}\begin{pmatrix}v_{1}\\ \\ v_{2}\end{pmatrix}=\begin{pmatrix}A&B\\ \\ C&D\end{pmatrix}\begin{pmatrix}v_{1}\\ \\ v_{2}\end{pmatrix}$, (2)
where, $q$ and $r$ are functions of $(x,t)$ whereas $A$, $B$, $C$ and $D$ depend on x,tx,\,t and $\zeta$, which may eventually depend on $t$. The field variables $x$ and $t$ can be regarded as space and time, respectively.
Equation (1) is a scattering problem while (2) is an evolution problem. From the cross differentiation and the integrability condition
$\partial_{t}\partial_{x}\begin{pmatrix}v_{1}\\ \\ v_{2}\end{pmatrix}=\partial_{x}\partial_{t}\begin{pmatrix}v_{1}\\ \\ v_{2}\end{pmatrix}$,
(1) and (2) give
$\begin{pmatrix}\partial_{x}A&\partial_{x}B\\[2.0pt] \partial_{x}C&\partial_{x}D\end{pmatrix}-\begin{pmatrix}0&\partial_{t}q\\[2.0pt] \partial_{t}r&0\end{pmatrix}=i\partial_{t}\zeta\begin{pmatrix}-1&0\\[2.0pt] 0&1\end{pmatrix}+\begin{pmatrix}qC-rB&q(D-A)-2i\zeta B\\[2.0pt] r(A-D)+2i\zeta C&rB-qC\end{pmatrix}$.
The $2\times 2$ matrix in (1) is traceless, as are all but the first in the above equation. If we impose that the second matrix in (2) be traceless, then all matrices we have so far will share this property. Without loss of generality, we may assume D=AD=-A.
@@
$\left\{\begin{array}[]{lcl}\partial_{x}A&=&qC-rB,\\ \\ \partial_{x}B+2i\zeta B&=&\partial_{t}q-2Aq,\\ \\ \partial_{x}C-2i\zeta C&=&\partial_{t}r+2Ar.\end{array}\right$. (3)
The reader may be wondering: What is the relevance of (3)? This is a very good question.
In [1] it was shown that a number of very important models, such as the KdV and the sine-Gordon equations, can be obtained from (3) depending on the choices of the involved functions. Moreover, equations arising from the compatibility condition (3) are said to be integrable (à la AKNS).
Let us give an example illustrating how an equation can be obtained from (3).
Example $1.$
@@
Then (3) is equivalent to the sine-Gordon (sG) equation
$u_{tx}=\sin{u}.$ (4)
It is worth mentioning that the sG equation had been discovered more than a century before [1] brought it to light, in a purely geometric context. In fact, the sG plays a vital role in the context of the geometry of surfaces with constant and negative Gaussian curvature [37, Chapter 1] and from its solutions it is possible to construct surfaces like those shown in Figure 1.
Figure 1: Surfaces of Gaussian curvature ${\cal K}=-1$. In [37, Chapter 1] it is shown how the above surfaces can be obtained from the solutions of the sine-Gordon equation (4).
Let us now introduce a geometric view on (1)–(2). By $\{dx,dt\}$ we denote the usual dual basis of $\mathbb{R}^{2}$. Consider eight one-forms $\Omega_{ij},\,\Psi_{ij}$, $1\leq i,j\leq 2$ and define the $2\times 2$ one-form matrices $\Omega=(\Omega_{ij})$, $\Psi=(\Psi_{ij})$. If $d$ denotes the usual exterior derivative, then we have
$d\Omega=(d\Omega_{ij}),$
@@
$\Omega=\begin{pmatrix}-i\zeta dx+Adt&qdx+Bdt\\ \\ rdx+Cdt&i\zeta dx-Adt\end{pmatrix}$ (5)
@@
$dv=\Omega v,$ (6)
where $v=(v_{1}\quad v_{2})^{t}$ (the superscript $t$ means the usual matrix transposition). From (5)–(6) we obtain
@@
$d\Omega=\Omega\wedge\Omega.$ (7)
@@
then (7) gives the equation
$\partial_{t}X+\partial_{x}T+[X,T]=0.$ (8)
Equation (8) is better known as the zero curvature representation, and it reveals to us that (1)–(2) can be seen as a geometric object.
Equations (6) is a different, but equivalent, way to write (1)–(2), whereas (7) implies to (3).
The remaining part of the paper is as follows: in Section 2, we revisit the main aspects of the works by Sasaki [38] and Chern and Tenenblat [5], connecting equations obtained from the AKNS method and the geometry of abstract 2-manifolds with constant and negative Gaussian curvature. In Section 3, we discuss some recent results from the author [29, 19, 16, 18, 14] that connect the theory developed by Chern and Tenenblat with problems concerning not only equations, but equations subject to some condition (initial data). In some parts, the presentation may closely follow the original references.
@@ 2 Equations describing pseudospherical surfaces
The study of relations between solutions of equations and the curvature of surfaces dates back to 1862, when, according to [37, Introduction], Edmond Bour discovered (4). Soon after, Bonnet and Enneper independently rediscovered the same equation in 1867 and 1868, respectively.
Although the works by Bour, Bonnet, and Enneper only considered the sine-Gordon equation, they can be seen as the light dawn of what today is better known as equations describing pseudospherical surfaces or PSS equations (henceforth abbreviated as PSS, also used in the plural).
After the mentioned works, Sasaki’s observation, connecting the AKNS system [1] with the structure equations for a pseudospherical surface, and later the fundamental work by Chern and Tenenblat, where the concept of PSS equations [5] was established, made connections between certain partial differential equations and smooth two-dimensional Riemannian manifolds of constant Gaussian curvature. Their contribution gained particular significance following a notable insight by Sasaki, who had earlier demonstrated that solutions to specific integrable equations—those arising from the AKNS formalism [1]—are intimately related to the intrinsic geometry of surfaces with constant negative Gaussian curvature ${\cal K}=-1$ [38].
By definition, a PSS is a two-dimensional Riemannian manifold whose Gaussian curvature remains constant and negative. According to Hilbert’s theorem [12, Section 5-11], such manifolds cannot be isometrically immersed in $\mathbb{R}^{3}$ as complete surfaces, due to their negative curvature—see also [30, page 439] and [27]. For this reason, they are frequently referred to as abstract surfaces in the literature.
${\cal E}(x,t,u,u_{(1)},\cdots,u_{(n)})=0.$ (9)
@@
A differential equation (9) is said to describe pseudospherical surfaces, or it is said to be of pseudospherical type, if it is a necessary and sufficient condition for the existence of differentiable functions $f_{ij}$, $1\leq i,j\leq 3$, such that the forms
$\omega_{i}=f_{i1}dx+f_{i2}dt,\quad 1\leq i\leq 3,$ (10)
satisfy the structure equations of a pseudospherical surface
$d\omega_{1}=\omega_{3}\wedge\omega_{2},\quad d\omega_{2}=\omega_{1}\wedge\omega_{3},\quad d\omega_{3}=\omega_{1}\wedge\omega_{2}.$ (11)
Sasaki noticed in his work [38] that by defining one-forms $\omega_{1}$, $\omega_{2}$ and $\omega_{3}$ suitably from (1) and (2), one could obtain the one-form matrix
$\Omega=\frac{1}{2}\begin{pmatrix}\omega_{2}&\omega_{1}-\omega_{3}\\ \omega_{1}+\omega_{3}&-\omega_{2}\end{pmatrix}$, (12)
and we can define
$\Sigma:=d\Omega-\Omega\wedge\Omega.$ (13)
Then $\Sigma=0$ if and only if (11) holds. As a result, if (9) is the compatibility condition of (3), then its solutions satisfy (11), that is, the structure equations for a two-dimensional Riemannian manifold with first fundamental form, or metric, given by
$I=\omega_{1}^{2}+\omega_{2}^{2}.$ (14)
These foundational results laid the groundwork for a new intersection between surface theory and the analysis of PDEs. Since then, this perspective has shaped a substantial body of literature, including [2, 40, 39, 4, 29, 19, 32, 33, 35, 34, 36, 14], among many others. A summary of Tenenblat’s works can be found in [41, Chapter 1], whereas [32] deals with geometric aspects of integrability. A nice survey of PSS equations can be found in the paper by Reyes [36].
At least in its roots, the works after [5] dealing with PSS equations were mostly concerned with integrable equations in the AKNS sense. As such, it was quite natural the presence of a fundamental parameter (in the sense it could not be removed after a Gauge transformation) in the one-forms (10).
From the point of view of AKNS or Lax integrability (see [13] for further details), the existence of such a parameter, known as the spectral parameter due to the fact that it appears in an eigenvalue problem, is somewhat related to the existence of infinitely many linearly independent integrals of motion or conserved quantities. This fact was first shown for the KdV by Miura, Gardner and Kruskal [28] and then explained in terms of a pair of operators by Lax [26] and named after him.
From a geometric perspective, the existence of the spectral parameter in the one-forms determined by the solutions of PSS equations implies the existence of a one-parameter family of PSS. This fact leads us to the notion of geometrically integrable equations, that are those equations whose solutions describe non-trivial one-parameter families of surfaces. To the best of my knowledge, this characterization was first stated by Reyes, see [32, Definition 2], although it dates back to the original work by Chern and Tenenblat, see the comment after Definition 2 in [32].
Despite its importance in the field of integrable systems, soon after the works [38, 5], the spectral parameter lost importance and the investigation of PSS equations shifted to problems of diverse nature over time, evolving to classification problems and more geometric questions, such as the problems of how immersion of surfaces may depend on jets of finite order, e.g, see [22, 23, 24].
In practical terms, given a triad of one-forms $\omega_{1}$, $\omega_{2}$ and $\omega_{3}$, and an equation (9), we can check if they describe a PSS surface in the following way: let us define a matrix of one-forms $\Omega$ by (12). If, when restricted to the manifold determined by the solutions [^2] The one-forms involved are, in fact, pullbacks of the one-forms by holonomic sections of jets. For a more geometrically rigorous formulation of these concepts, see [31, Section II], see also Remark 3.1. of (9), the matrix $\Sigma$ vanishes, we say that (9) is a PSS equation, and the triad $\{\omega_{1},\,\omega_{2},\,\omega_{3}\}$ satisfies the structure equations (11) of a PSS with Gaussian curvature ${\cal K}=-1$. As a result, any open, simply connected set $U$ where the one-forms $\omega_{1}$ and $\omega_{2}$ are defined and $\omega_{1}\wedge\omega_{2}\neq 0$ everywhere, is endowed with a PSS structure. We can naively see this condition as a requirement for $\{\omega_{1},\omega_{2}\}$ to be a linearly independent set and thus be the basis of a plane.
@@
Assume that (9) is an equation describing a PSS and let $u$ be a solution of it. We say that $u$ is a generic solution if the one-forms (10) satisfy the condition $\omega_{1}\wedge\omega_{2}\neq 0$.
@@
$\left\{\begin{array}[]{lcl}\omega_{1}&=&{\frac{1}{\eta}\sin{u}dt,}\\ \\ \omega_{2}&=&{\eta dx+\frac{1}{\eta}\cos{u}dt},\\ \\ \omega_{3}&=&u_{x}dx.\end{array}\right$. (15)
@@
$\left\{\begin{array}[]{lcl}d\omega_{1}-\omega_{3}\wedge\omega_{2}&\equiv&{0}\\ \\ d\omega_{2}-\omega_{1}\wedge\omega_{3}&\equiv&{0},\\ \\ d\omega_{3}-\omega_{1}\wedge\omega_{2}&=&-(u_{tx}-\sin{u})dx\wedge dt.\end{array}\right$.
Thus, on the solutions of the sG equation (4), the forms (15) satisfy (11). Since
@@
then $u$ will be a generic solution for (4) provided that $u\notin\{k\pi,\,\,\,k\in\mathbb{Z}\}=:P$. In particular, if $u$ is a non-constant solution, then it is non-generic, and its domain is ensured with a PSS structure in a simply connected neighborhood of $(x,t)$ whenever $u(x,t)\notin P$.
@@
$\begin{array}[]{lcl}\omega_{1}&=&{\Big(\frac{\lambda}{2}+\frac{1}{2\lambda}-m\Big)dx+\Big(um+\frac{\lambda}{2}u-\frac{u}{2\lambda}-\frac{1}{2}-\frac{\lambda^{2}}{2}\Big)dt},\\ \\ \omega_{2}&=&-u_{x}dt,\\ \\ \omega_{3}&=&{\Big(m+\frac{1}{2\lambda}-\frac{\lambda}{2}\Big)dx+\Big(\frac{\lambda^{2}}{2}-\frac{1}{2}-\frac{u}{2\lambda}-\frac{\lambda}{2}u-um\Big)dt}.\end{array}$ (16)
@@
$\begin{array}[]{lcl}d\omega_{1}-\omega_{3}\wedge\omega_{2}&=&\Big(u_{t}-u_{txx}+3uu_{x}-2u_{x}u_{xx}-uu_{xxx}\Big)dx\wedge dt,\\ \\ d\omega_{2}-\omega_{1}\wedge\omega_{3}&=&0,\\ \\ d\omega_{3}-\omega_{1}\wedge\omega_{2}&=&-\Big(u_{t}-u_{txx}+3uu_{x}-2u_{x}u_{xx}-uu_{xxx}\Big)dx\wedge dt,\end{array}$ (17)
meaning that on the generic solutions of the CH equation
$u_{t}-u_{txx}+3uu_{x}=2u_{x}u_{xx}+uu_{xxx},$ (18)
or its equivalent form
$m_{t}+2um_{x}+u_{x}m=0,$ (19)
the one-forms (16) define a two-dimensional Riemannian manifold with first fundamental form
$\begin{array}[]{lcl}I&=&{\Big[m^{2}-\Big(\lambda+\frac{1}{\lambda}\Big)m+\frac{1}{4}\Big(\lambda+\frac{1}{\lambda}\Big)^{2}\Big]dx^{2}}+\\ \\ &&{\Big[-um^{2}+\frac{1}{\lambda}um+\frac{1}{4}\Big(\lambda^{2}-\frac{1}{\lambda^{2}}\Big)u+\lambda\Big(\frac{\lambda}{2}+\frac{1}{2\lambda}\Big)m-\lambda\Big(\frac{\lambda}{2}+\frac{1}{2\lambda}\Big)^{2}\Big]dxdt}+\\ \\ &&{\Big[um^{2}+u^{2}m\Big(\lambda-\frac{1}{\lambda}\Big)+u_{x}^{2}-\lambda\Big(\lambda+\frac{1}{\lambda}\Big)um+\frac{1}{4}\Big(\lambda+\frac{1}{\lambda}\Big)^{2}u^{2}-\frac{\lambda}{2}\Big(\lambda+\frac{1}{\lambda}\Big)^{2}}\\ \\ &&{+\frac{\lambda^{2}}{4}\Big(\lambda+\frac{1}{\lambda}\Big)^{2}\Big]dt^{2}}.\end{array}$
@@
$m=\frac{\lambda}{2}+\frac{1}{2\lambda}$ (20)
@@
Remark 2.5 (Non-uniqueness of the one-forms).
It is possible that a PSS equation be the compatibility condition for more than one triad of one forms satisfying the structure equations (11). To illustrate this observation, consider the one-forms
$\begin{array}[]{lcl}\omega_{1}&=&{\Big(u_{xx}-u-\beta+\frac{\beta}{\eta^{2}}-\frac{1}{\eta^{2}}\Big)dx}\\ \\ &&+{\Big(-\frac{\beta}{\eta}u_{x}-\frac{\beta}{\eta^{2}}+u^{2}-1+\beta u+\frac{u_{x}}{\eta}+\frac{1}{\eta^{2}}-uu_{xx}\Big)dt}\\ \\ \omega_{2}&=&{\eta dx+\Big(-\frac{\beta}{\eta}-\eta u+\frac{1}{\eta}+u_{x}\Big)dt},\\ \\ \omega_{3}&=&{\Big(u_{xx}-u+1\Big)dx+\Big(\frac{\beta}{\eta^{2}}u+u^{2}=uu_{xx}+\frac{1}{\eta^{2}}+\frac{u_{x}}{\eta}-\frac{u}{\eta^{2}}-u-\frac{\beta}{\eta^{2}}-\frac{\beta}{\eta}u_{x}\Big)dt},\end{array}$
where $\eta^{4}-\eta^{2}+\beta^{2}\eta^{2}=\beta^{2}+1-2\beta$. In [33, Theorem 1] it was shown that if we substitute the one-forms above into (11), we then obtain the CH equation (18) as the compatibility condition. Compare with the representation in Example 2.4.
Remark 2.6 (Geometric integrability and AKNS integrability are not equivalent).